### v1

# Learning from Emotions, Demographic Information and Implicit User Feedback in Task-Oriented Document-Grounded Dialogues

Abstract
The success of task-oriented and document-grounded dialogue systems depends on users accepting and enjoying using them. To achieve this, recently published work in the field of Human-Computer Interaction suggests that the combination of considering demographic information, user emotions and learning from the implicit feedback in their utterances, is particularly important. However, these findings have not yet been transferred to the field of Natural Language Processing, where these data are primarily studied separately. Accordingly, no sufficiently annotated dataset is available. To address this gap, we introduce FEDI, the first English dialogue dataset for task-oriented document-grounded dialogues annotated with demographic information, user emotions and implicit feedback. Our experiments with FLAN-T5, GPT-2 and LLaMA-2 show that these data have the potential to improve task completion and the factual consistency of the generated responses and user acceptance111 Code and data is available in GitHub..
Dataset Source Type Demographics Emotions
Implicit
Feedback
#Dialogues
Avg. Num.
of Turns
Avg. Utt.
Length
Lexical
Diversity
FITS
Xu et al. (2023b)
Crowdsourced
Document-
Grounded
22k 7.1 15.0 52.8
Blenderbot 3x
Xu et al. (2023a)
Open-Domain 261k 11.3 14.2 47.3
SaferDialogues
Ung et al. (2022)
8k 2.5 14.8 53.3
EmotionLines
Hsu et al. (2018)
1k 7.3 7.8 68.5
EmpatheticDialogues
Rashkin et al. (2019)
25k 4.3 13.7 64.2
SODA
Kim et al. (2023)
LLM- Generated Open-Domain 1.5M 7.6 16.1 68.0
PersonaChatGen
Lee et al. (2022)
1.6k 16.0 9.5 56.7
FEDI
LLM-
Generated
Task-Oriented
Document-
Grounded
8.8k 7.6 16.8 62.1
Table 1: Comparison of FEDI to other datasets that provide related annotations. FEDI is comparable with other LLM-generated datasets in terms of avg. turn and utterance length, and has a higher lexical diversity (based on MTLD McCarthy and Jarvis (2010) [^3] We used the Python package lexical-diversity v0.1.1 for calculation (last accessed 04 January 2024).) than many of the crowdsourced datasets.
1 Introduction
Figure 1: A feedback dialogue from FEDI. User emotion and implicit feedback annotations (generation error and user reaction types) are beneath the utterances.
Task-oriented and document-grounded dialogue systems, such as personal assistants or customer service agents, often implement sensitive tasks, which is why user acceptance and trust are important success criteria. Recently published work in the field of Human-Computer Interaction (HCI) suggests that considering demographic information (such as age, occupation or language style), interpreting user emotions, and learning from implicit user feedback, such as a correction or question in response to a generation error in the preceding system utterance, are particularly important because they often interact and correlate with each other Pelau et al. (2021); Araujo (2018); Zamora (2017); Ciechanowski et al. (2019). For example, in the second utterance of Figure 1 the system misinterprets Claudia’s question, resulting in an incorrect response. This has a direct impact on her emotional state and she now asks the system for clarification (implicit feedback). After the system has generated a more suitable response, she is satisfied, her emotional state changes again and she asks the system another question.
As of today, these findings have not yet been transferred to research in the field of Natural Language Processing (NLP), where these data are usually considered separately from each other. As a result, a large number of works are available that either focus on learning from demographic information Lee et al. (2022); Zhang et al. (2018), user emotions Firdaus et al. (2020); Rashkin et al. (2019); Hsu et al. (2018); Hwang et al. (2023) or implicit user feedback Hancock et al. (2019); Veron et al. (2021); Xu et al. (2023a, b); Ung et al. (2022), but none that considers them together, which is a limitation for future research Chaves and Gerosa (2021); Minjin Rheu and Huh-Yoo (2021).
In this work, we address this gap by (a) providing FEDI, the first English dialogue dataset annotated with implicit user Feedback, Emotions and Demographic Information, and (b) investigating the effect of including these data as part of the input on the performance of three different state-of-the-art language generation models, i.e., FLAN-T5 Chung et al. (2022), GPT-2 Radford et al. (2019) and LLaMA-2 Touvron et al. (2023). FEDI consists of overall 8k task-oriented document-grounded dialogues, 6k of which are annotated with implicit user feedback.
To create FEDI in a cost and resource efficient way, we use GPT-3.5-Turbo [^4] OpenAI GPT-3.5 Model Page (last accessed on 02 January 2024). The model is based on Ouyang et al. (2022). The data was generated between March and June 2023. to generate and annotate the training and validation data, and recruit humans to assess its quality and to collect a separate set of test dialogues, due to the potential limitations of synthetic data Ji et al. (2023); Zhang et al. (2023); Malaviya et al. (2023). We also use GPT-3.5-Turbo for the generation of implicit user feedback. In contrast to previous works which usually focus on feedback specific to their use case Xu et al. (2023b); Ung et al. (2022); Hancock et al. (2019), we use the taxonomies proposed by Petrak et al. (2023) for this, since they cover a variety of different feedback situations, such as feedback on ignored user requests, knowledge gaps or harmful social behavior. In summary, we provide these contributions:
  1. 1.
    FEDI, the first task-oriented document-grounded dialogue dataset for learning from demographic information, user emotions and implicit user feedback.
  2. 2.
    New experimental insights showing that learning from demographic information, user emotions and implicit feedback has a positive impact on task completion and factual consistency of generated responses in task-oriented knowledge-grounded dialogues.
  3. 3.
    A new framework for generating and annotating high-quality task-oriented knowledge-grounded feedback-annotated dialogue data.
FEDI is comparable to other related human-bot and LLM-generated datasets in terms of size, lexical diversity and dialogue length (see Table 3). In our analysis, we provide insights into the quality of generated annotations, especially in comparison to the test data collected by humans.
2 Related Work
Demographic information, user emotions and implicit user feedback are usually considered separately in the field of NLP. However, available works show that considering demographic information on dialogue-level and as part of the dialogue context has a positive impact on the user experience Lee et al. (2022); Zhang et al. (2018); Hwang et al. (2023); Siddique et al. (2022); Luo et al. (2019). This is similar for user emotions, which are usually considered on utterance-level and detected by separate modules, such as emotion detectors Firdaus et al. (2020); Rashkin et al. (2019); Hsu et al. (2018). To improve upon implicit user feedback usually requires to train the model with the feedback data Ung et al. (2022); Xu et al. (2023a); Veron et al. (2021). In this regard, continual learning has shown to be very promising Xu et al. (2023b); Wang et al. (2019); Hancock et al. (2019). Table 3 gives a comparison of the datasets resulting from these works. It shows that (1) no dataset is available for task-oriented dialogues, and (2) most of the available datasets are the result of resource-intensive crowdsourcing efforts, often leading to datasets of varying quality due to, e.g., methodological artifacts or annotator biases Yang et al. (2023); Parmar et al. (2023); Thorn Jakobsen et al. (2022); Prabhakaran et al. (2021). As an alternative, recent works suggest synthetic data generation using large language models as a more efficient approach to generate high-quality dialogue data Kim et al. (2023); Li et al. (2023); Lee et al. (2022), despite their limitations, such as a tendency to generate hallucinated or harmful output Ji et al. (2023); Zhang et al. (2023); Malaviya et al. (2023).
In this work, we use GPT-3.5-Turbo to generate a task-oriented document-grounded dialogue dataset annotated with demographic information, user emotions and implicit feedback to investigate the impact of such information on the performance of three state-of-the-art language generation models. Due to the potential limitations of synthetic data, we recruit human annotators for quality assessment, curation, and to collect a separate set of test dialogs.
Figure 2: Overview of our framework for generating and annotating dialogues. We distinguish feedback-free and feedback dialogues. Feedback scenarios and error and user reactions types (marked with a *) are specific to feedback dialogues. We provide GPT-3.5-Turbo with all the information from the box on the left side to (1) generate the dialogue in Step 1 and (2) generate the slot and emotion annotations in Step 2.
3 FEDI
FEDI covers four use cases for task-oriented document-grounded dialogue systems from three domains, including post office services, receptionist services and customer services in the insurance domain. For post office services, we include (1) customer support for parcel shipping, i.e., guiding them through the process of parcel shipping from choosing the right shipping box to informing them about the approximate delivery time, and (2) topping up a prepaid SIM card. For reception and customer services in the insurance domain, we include one use case each, i.e., access control (the reception and registration of new visitors in office buildings) and question answering (in the context of financial topics and pet, health and heritage insurance [^5] We use the policies of POSTE Italiane as a reference for this, which are also publicly available in English (last accessed 04 January 2024).). The question answering dialogues are additionally annotated with the documents that provide the knowledge required for response generation. Appendix A describes the tasks in more detail, including slots, intents and examples.
Demographic Information
We consider gender, age, occupation, name, and language style as demographic information in this work. Overall, we distinguish 12 different language styles, such as formal, dialect and jargon, five demographic cohorts, ranging from Boomers (born between 1952 and 1962) to Generation Alpha (born between 2007 and 2016), and a variety of $1,155$ occupations and $1,000$ names, sampled from The Gazette [^6] Available in GitHub (last accessed on 04 January 2024).. We provide more details in Appendix B.
User Emotions
For user emotions, we use the taxonomy from EmotionLines Hsu et al. (2018), which covers seven different emotions, including Neutral, Joy (which we refer to as Happiness), Sadness, Surprise, Fear, Anger, and Disgust. We extend this list with four emotion types we found to be relevant in related work Kim et al. (2023); Rashkin et al. (2019), including Confusion, Curiosity, Frustration, and Stress. We consider Confusion, Frustration, Fear, Sadness, Disgust, Stress, and Anger as negative emotions.
Implicit User Feedback
For the generation and annotation of implicit user feedback, we use the user reaction type taxonomy proposed by Petrak et al. (2023), which distinguishes five user reaction types in response to generation errors in preceding system utterances, including (1) Ignore and Continue, (2) Repeat or Rephrase, (3) Make Aware With Correction, (4) Make Aware Without Correction, and (5) Ask for Clarification. (2), (3) and (5) describe user reactions that contain implicit user feedback. For generation errors in system utterances, they also propose an error taxonomy of ten types, nine of which are relevant for task-oriented document-grounded dialogues, such as Ignore Request, Attribute Error, Factually Incorrect, or Lack of Sociality. We only consider these error types for generating implicit user feedback. Definitions, further details and examples can be found in Appendix B.
Problem Formulation
We define a dialogue as a set of multiple turns $T$. Each turn consists of two utterances, a user utterance $U_{t}$ and a system utterance $S_{t}$. Given the dialogue context $C=[T_{0},...,T_{t-1}]$, and additional information $K$, the task is to predict the user intent $I_{t}$, generate belief state $B_{t}$ and system utterance $S_{t}$:
$(I_{t},B_{t},S_{t})=\text{generate}(K,C,U_{t})$ (1)
Depending on whether knowledge from a document $D_{t}$ is required to generate $S_{t}$ or the user emotion $E_{t}$, demographic information $DI$, generation error $GE_{t}$, or implicit user feedback $F_{t}$ should be considered, $K=\{D_{t},DI,E_{t},GE_{t},F_{t}\}$. $DI$ includes the user’s gender, age range, occupation, name, and language style. Belief state $B_{t}$ includes the slot values inferred from the dialogue context $C$, which may be used to query knowledge from an external information retrieval system Chen et al. (2022); Hosseini-Asl et al. (2020), such as the document $D_{t}$ in the case of customer service or registration information in the case of access control. We do not use such an additional retrieval system in this work and provide all the information required to generate the response $S_{t}$ as part of the additional information $K$ in the input sequence.
4 Framework for Generating and Annotating Dialogues
Figure 2 gives an overview of our framework for generating and annotating dialogues using GPT-3.5-Turbo. In general, we distinguish feedback-free and feedback dialogues, i.e., dialogues that provide annotations for implicit user feedback. However, the procedure for dialogue and annotation generation is in general the same for both (Section 4.1). Section 4.2 describes the steps specific to feedback dialogues. For each step that involves GPT-3.5-Turbo, we require the model to return the results in a predefined JSON scheme. If in one step the generation does not match this requirement, does not provide all required fields or contains invalid encoding, the whole dialogue is discarded. We provide more details, including the instructions used in this procedure, in Appendix C.
4.1 General Approach to Dialogue Generation
For dialogue generation, we provide GPT-3.5-Turbo with the randomly sampled demographic information for the user, a task description, and the role of the starting actor, i.e., user or system. As indicated by the illustration on the left side of Figure 2, a task description describes the flow of events and information which needs to be conveyed by each role to fulfill the task. In the case of question answering, the task description also includes a randomly sampled list of documents from the respective topic. Similar to PersonaChatGen Lee et al. (2022), we instruct the model to use the task description and the demographic information to generate a background story to guide the conversation, such as depicted in the center of Figure 2. We also instruct the model to return the utterance-level annotations for intents and limit the dialogue to 13 turns, since we found that longer dialogues tend to deviate from the task description. For background stories, we limit the length to five sentences, since we found that longer background stories tend to be distracting.
Annotation Generation
The generation of annotations is the second step in our generation framework [^7] We also tried to reduce API calls by combining dialogue and annotation generation but found that this does not produce reliable results. (see Figure 2). For slot annotations, we provide GPT-3.5-Turbo with the generated dialogue and a list of all slots defined in the task description, possible values and examples. We also instruct the model to only assign and copy values from the dialogue (to prevent hallucinations) and to return the annotations on utterance-level. For emotion annotations, we instruct the model to predict the emotion for each user utterance in the dialogue, given the dialogue and our emotion taxonomy.
4.2 Feedback Dialogues
Feedback Scenarios
For each feedback dialogue, we first use GPT-3.5-Turbo in a prior step to generate three feedback scenarios that are then used as an additional source for dialogue generation. A feedback scenario describes a generation error and the following implicit user feedback (the center of Figure 2 shows an example) [^8] We generate all feedback scenarios for a dialogue at once, using the same API call.. For this, we provide the model with the task description and a list of randomly sampled error and user reaction types. To ensure coherence, we add the constraint that feedback scenarios must be non-mutually exclusive and together form a story in the context of the task description. We limit the number of feedback scenarios per dialogue to not exceed the length of 13 turns per dialogue and to leave room for the start and end of the conversation.
Feedback Dialogue Generation
For feedback dialogue generation, we instruct GPT-3.5-Turbo to consider each feedback scenario in three utterances in the generated dialogue, in addition to the constraints described in Section 4.1: The system utterance with the generation error, a subsequent user utterance that reflects the user reaction, and a following system utterance that addresses the user reaction. Next, we generate three additional versions of the same dialogue, each resolving one of the feedback scenarios. Figure 3 illustrates this procedure.
Figure 3: During the generation of feedback dialogues, we subsequently resolve feedback scenarios in the original feedback dialogues.
We first mask the system utterance with generation error and generate a replacement using the task description and the preceding dialogue context. Next, we drop the following two utterances, since they are directly related to the original generation error. This way, the conversation continues with the next regular user utterance. We continue the process until all feedback turns have been resolved as in Version 4.
5 FEDI Analysis
FEDI consists of 8,852 dialogues, divided into 1,988 feedback-free dialogues, including 326 test dialogues, and 6,864 feedback dialogues (1,716 in four versions, each with one feedback scenario less per dialogue). The test dialogues were collected human-human by eight computer science students in overall 136 paid working hours (see Appendix D for the procedure and details on the hiring process and salary). In the following, we focus on the completeness of generated slot and intent annotations, the distribution of user emotions and the feedback scenarios represented in the dialogues. We provide additional statistical analysis in Appendix E, including split sizes and the distribution of demographic information.
Slot and Intent Annotations
Table 10 shows the ratio of dialogues for which intent and slot annotations were successful, i.e., dialogues that provide all annotations for intent and required slot values.
Task
Feedback-Free
Dialogues
Feedback Dialogues
Gen. Test Version 1 Version 2 Version 3 Version 4
Parcel
Shipping
0.87 0.51 0.74 0.72 0.70 0.70
Top Up
SIM Card
0.87 0.51 0.74 0.72 0.71 0.69
Access
Control
0.86 0.68 0.82 0.83 0.84 0.84
Question
Answering
0.99 0.87 0.73 0.99 0.99 0.99
Table 2: The table shows the ratio of dialogues that are complete in the sense that they are annotated with all intent and slot values [^10] Hallucinated slot values, i.e. slot annotations that do not occur in the respective utterance, are counted as missing..
For the feedback-free dialogues, we distinguish between generated dialogues (Gen.) and test dialogues (Test, collected by humans). We observe large differences between (1) question answering and the other tasks and (2) generated dialogues and the test dialogues collected by humans. We found that this is mostly due to variations in the slot annotations. While the slot annotation scheme for question answering is rather simple (see Appendix A), this is different for the other tasks where slots often depend on the the background story. For example, in the case of parcel shipping, if the user already has a shipping box and just requires information on the shipping procedure, details about available shipping box types are negligible. While human annotators take this into account and occasionally omit slots that are not required based on dialogue motivation, GPT-3.5-Turbo follows our instructions more closely, which include all slots as part of the task description. For feedback dialogues, we observe that the generated corrections not always address the missing information required by the task description. We provide more analysis as part of our human curation study in Section 6.
Emotion Annotations
Figure 4 shows the distribution of the five most common emotions observed in user utterances from both the feedback-free and feedback dialogues [^11] We do not distinguish between generated and test dialogues here. We also leave out the neutral emotion as it is in general the most frequently observed emotion (40.5% of all annotated emotions)..
Figure 4: Ratio of the most commonly observed user emotions in FEDI (excluding the Neutral emotion).
As expected, negative emotions are more common in feedback dialogues. For Curiosity, we found that the polarity depends on the dialogue context, e.g. whether the previous system utterance successfully addressed the user’s request. It is an emotion that can be either positive or negative, thus it is frequently observed in both dialogues types. Happiness in feedback dialogues is mostly observed as a reaction to system utterances that implement user feedback.
Feedback Scenarios
Figure 5 shows the distribution of user reactions in relation to error types represented in the feedback scenarios of the feedback dialogues.
Figure 5: Distribution of user reactions in relation to error types represented in feedback scenarios.
The figure shows that our approach for generating feedback scenarios resulted in meaningful combinations of error and user reaction types. For example, Factually Incorrect is mostly addressed by Make Aware with Correction. Ignore and Continue are only rarely observed. The same applies to Unclear Intention and Attribute Error, which are mostly addressed by Ask for Clarification and Repeat or Rephrase. The latter one is also frequently observed in combination with Ignore Question and Ignore Expectation errors.
Experiment Matching Task Completion Quality
F1 BLEU BertScore Inform Success Intent Acc Slot Acc Toxicity
FLAN-T5 Feedback-Free FLAN-T5 45.0 20.0 88.3 86.7 85.9 54.8 60.9 0.02 52.7
 +Emotions 46.7 (+1.7) 21.0 (+1.0) 88.9 (+0.6) 83.9 (–2.8) 83.2 (–2.7) 61.2 (+6.4) 58.3 (–2.6) 0.02 57.5 (+4.8)
 +Demographics 43.2 (–1.8) 18.4 (–1.6) 87.7 (–0.6) 87.0 (+0.3) 86.0 (+0.1) 33.5 (–21.3) 29.3 (–31.6) 0.03 (+0.01) 54.5 (+1.8)
 +Emotions
 +Demographics
44.2 (–0.8) 19.1 (–0.9) 88.1 (–0.2) 85.3 (–1.4) 85.1 (–0.8) 43.9 (–10.9) 36.7 (–24.2) 0.02 56.4 (+3.7)
Feedback  +Generation Error 41.4 (–3.6) 19.8 (–0.2) 87.8 (–0.5) 96.8 (+10.1) 92.7 (+6.8) 72.5 (+17.7) 76.7 (+15.8) 0.02 56.9 (+4.2)
 +User Reaction 41.3 (–3.7) 19.3 –0.7 87.6 (–0.7) 96.6 (+9.9) 94.1 (+8.2) 69.0 (+14.2) 76.2 +15.3 0.02 56.3 (+4.1)
 +Generation Error
 +User Reaction
43.4 (–0.6) 22.1 (+2.1) 88.2 (–0.1) 96.9 (+10.2) 95.3 (+9.4) 83.5 (+28.7) 77.2 (+16.3) 0.02 60.2 (+7.5)
GPT-2 Feedback-Free GPT-2 34.9 10.4 87.1 88.3 81.6 78.7 69.6 0.02 28.1
 +Emotions 35.1 (+0.2) 10.4 87.1 84.1 (–4.2) 83.8 (+2.2) 75.4 (–3.3) 67.3 (–2.3) 0.02 26.7 (–1.4)
 +Demographics 34.6 (–0.3) 10.4 87.1 80.2 (–0.1) 80.2 (–1.4) 69.3 (–9.4) 57.5 (–12.1) 0.02 26.3 (–1.8)
 +Emotions
 +Demographics
36.0 (+1.1) 11.4 (+1.0) 87.3 (+0.2) 85.1 (–3.2) 84.8 (+3.2) 71.6 (–7.1) 66.7 (– 2.9) 0.02 29.2 (+1.1)
Feedback  +Generation Error 29.2 (–5.7) 8.0 (–2.4) 86.2 (–0.9) 92.4 (+4.1) 91.7 (+10.1) 84.3 (+5.6) 79.3 (–9.7) 0.02 30.9 (+2.8)
 +User Reaction 30.0 (–4.9) 8.3 (–2.1) 86.3 (–0.8) 98.9 (+10.6) 96.5 (+14.9) 83.0 (+4.3) 80.3 (+10.7) 0.02 32.3 (+4.2)
 +Generation Error
 +User Reaction
30.3 (–4.6) 9.7 (–0.7) 86.4 (–0.7) 94.7 (+6.4) 93.3 (+11.7) 88.0 (+9.3) 80.8 (+11.2 0.01 (–0.01) 35.5 (+7.4)
LLaMA-2 Feedback-Free LLaMA-2 29.3 7.1 86.1 85.9 81.2 37.6 39.2 0.02 28.3
 +Emotions 36.3 (+7.0) 14.9 (+7.8) 85.4 (–0.7) 89.3 (+3.4) 85.3 (+4.1) 40.2 (+2.6) 41.3 (+2.1) 0.01 (–0.01) 18.7 (–9.6)
 +Demographics 33.8 (+4.5) 4.5 (–2.6) 86.5 (+0.4) 85.6 (–0.3) 82.5 (+1.3) 37.1 (–0.5) 40.1 (+0.9) 0.02 21.3 (–7.0)
 +Emotions
 +Demographics
28.8 (–0.5) 5.6 (–1.5) 81.3 (–4.8) 86.7 (+1.8) 87.9 (+6.7) 41.4 (+3.8) 39.6 (+0.4) 0.03 (+0.01) 20.6 (–7.7)
Feedback  +Generation Error 24.1 (–5.2) 7.9 (+0.8) 77.4 (–8.7) 93.1 (+7.2) 95.7 (+14.5) 54.8 (+17.2) 59.6 (+20.4) 0.01 (–0.01) 29.1 (+0.8)
 +User Reaction 24.5 (–4.8) 6.9 (–0.2) 78.8 (–7.3) 94.9 (+9.0) 93.2 (+12.0) 63.5 (+25.9) 70.1 (+30.9) 0.02 27.1 (–1.2)
 +Generation Error
 +User Reaction
25.0 (–4.3) 9.2 (+1.9) 80.1 (–6.0) 82.4 (–3.5) 83.6 (+2.4) 46.3 (+8.7) 47.2 (+8.0) 0.03 (+0.01) 33.5 (+5.0)
Table 3: Results of our experiments. We use the baseline models as deltas, i.e., the pretrained models finetuned on the FEDI feedback-free dialogues. The models with the greatest improvements are printed in bold.
6 Human Curation Study
To assess the quality of our generated dialogues, we used INCEpTION Klie et al. (2018) as a platform and asked two participants from our test data collection to assess and curate the intent, slot and emotion annotations in 480 feedback-free dialogues and the error and user reaction type annotations in 380 feedback dialogues. To measure the agreement between both annotators (IAA), we use Krippendorff’s Alpha Krippendorff (2006) as provided in the INCEpTION platform.
Annotation Type Missing Changed IAA
Feedback-Free Dialogues Intent 6% 35% 0.90
Slots 56% 19% 0.83
Emotion 2% 81% 0.91
Feedback Dialogues Error Type 16% 36% 0.97
User Reaction Type 16% 34% 0.89
Table 4: The ratio of dialogues with at least one missing or changed annotation in our human curation study.
Table 4 shows the results [^12] Overall, 26 dialogues were reported as off-topic and are not considered in these results.. Overall, the ratio of dialogues with at least one missing annotation is low, except for slot annotations. We found that most of them are parcel shipping dialogues, which has a comparatively complex annotation scheme (see Appendix A). We attribute this to our observation on slot and intent annotations from Section 5, i.e., the (occassional) dependence on the background story. A detailed analysis revealed that an average of 1.8 annotations were added to the dialogues, most of them (36%) were slot annotations. For the dialogues with at least one changed annotation, we found that in many of these cases placeholders, e.g., the slot name put in brackets ([shipping_box_name]), were used instead of the slot values from the dialogues (reported by the students). Emotion is the most frequently changed annotation type (on average 2.09 times per affected dialogue), with the originally-annotated emotion often being Neutral.
7 Experiments and Results
To investigate the impact of combining demographic information with user emotion and implicit feedback, we conduct experiments using three models of different architecture and pretraining approaches, including FLAN-T5 Chung et al. (2022) (780M), GPT-2 Radford et al. (2019) (780M) and LLaMA-2 Touvron et al. (2023) (7B) [^13] The model weights for FLAN-T5 and GPT-2 are available in the Huggingface Model Hub (last accessed 04 January 2024). Access to the weights for LLaMA-2 must be requested from Meta AI (last acessed 26 December 2023).. For LLaMA-2, we only finetune the LoRA Hu et al. (2022) weights in our experiments. Table 3 shows the results achieved in the test data averaged over three runs.
Experimental Setup
We first finetune the pretrained models to our scenario using the feedback-free dialogues (Feedback-Free in Table 3) and included the demographic information (+Demographics) and user emotions (+Emotions) as part of the input sequences. We then use the best performing feedback-free models (the bold ones) for experiments using the feedback dialogues (Feedback). We provide more details, including hyperparameters and input sequences, in Appendix F.
Evaluation Metrics
We use F1-Score (calculated based on overlapping tokens in target and prediction), BLEU(-n) Papineni et al. (2002) and BertScore Zhang et al. (2020) to evaluate the concordance (Matching) of the generated system utterances with the target values. To evaluate for task completion, we use Inform and Success as proposed by Budzianowski et al. (2018) and additionally measure the correctness of the predicted intents (Intent Accuracy) and slot values (Slot Accuracy). To measure the toxicity in the generated responses, we use Perspective API. To measure the factual consistency in the case of question answering, we use $\text{Q}^{2}$ Honovich et al. (2021) [^14] For BLEU and BertScore, we use the implementation from the HuggingFace evaluation library v0.4.1 (last accessed 26 December 2023) and with $n=4$ for BLEU. For Inform and Success, we use the implementation from Nekvinda and Dušek (2021) as a reference (last accessed 28 December 2023). For $Q^{2}$, we use the reference implementation which is available in GitHub (last accessed 29 December 2023). Perspective API is a free-to-use service provided by Google and Jigsaw. Model and training details can be found here (last accessed 04 January 2024)..
Results
In general, we find that including user emotions (+Emotions) has a positive impact. This is particularly significant in the case of LLaMA-2 Touvron et al. (2023), where we used instructions for finetuning that provide additional context, i.e., a brief description of how the user emotion should be considered (see Appendix F for input sequences). In the feedback experiments, we observe great improvements in task completion and factual consistency of generated responses ($Q^{2}$ metric), which are both crucial for task-oriented document-grounded dialogues. For example, by including the generation error in the case of LLaMA-2 Touvron et al. (2023) or by combining the generation error with the user reaction in the case of FLAN-T5 Chung et al. (2022) and GPT-2 Radford et al. (2019). In general, we attribute these improvements to the additional context provided by the generation error and the user reaction, which can be interpreted as a negative example for a response in the specific dialogue context.
Human Evaluation
We conduct a small-scale human evaluation with two participants from our lab [^15] One of the authors and an intern of our lab who participated during his working hours.. We use the two best feedback-trained models from Table 3 (FLAN-T5 and GPT-2 with generation error and user reaction) and their feedback-free counterparts (FLAN-T5 and GPT-2 with user emotions) to generate responses for 50 randomly chosen samples from the FEDI test set. We asked the two annotators to rate the generated responses for human-likeness (naturalness), relevancy in the dialogue context (coherence), social acceptability (safety), factual consistency (with the target document in the case of answering questions), and engagement (whether they would use this model in practice). We use a likert scale from 1 (lowest rating) to 5 (highest rating) for each attribute and provide the annotators with the knowledge document, dialogue context, and generated response for this evaluation. Table 5 shows the results.
Experiment Naturalness Coherence Safety Engagement
Factual
Consistency
FLAN-T5
Feedback-Free
4.14 4.15 4.55 3.75 2.20
Feedback
4.30 (+0.16) 4.25 (+0.10) 4.59 (+0.09) 3.89 (+0.14) 2.24 (+0.04)
GPT-2
Feedback-Free
4.24 3.82 4.45 3.44 1.55
Feedback
4.42 (+0.18) 4.05 (+0.23) 4.46 (+0.01) 3.76 (+0.32) 1.58 (+0.03)
Table 5: Results of our human evaluation.
The reponses generated by the feedback-trained FLAN-T5 model are mostly rated higher by the participants. They also reported that the generated responses encourage for more user interaction, e.g., by requesting additional information or paying more attention to the user and their situation, and are in general more factual consistent. In the case of question answering, the generated responses are mostly summaries of the respective documents. According to the authors, the GPT-2 model trained only with emotions already produced very engaging answers, although they are not as coherent as the responses generated by FLAN-T5. This also affects factual consistency in the case of question answering, which is much lower for both GPT-2 models than for FLAN-T5.
8 Conclusion
In this work, we introduced FEDI, the first dataset for learning from demographic information, user emotions and implicit feedback in task-oriented document-grounded dialogues. FEDI consists of 8.8k dialogues generated and annotated using GPT-3.5-Turbo, 6k of which with annotations for implicit user feedback. For testing, we additionally provide a set of dialogues collected by humans. Our analysis shows the effectiveness of our generation framework and that FEDI is comparable to other related datasets in terms of size, utterance length, turn length, and lexical diversity. Our experiments with FLAN-T5, GPT-2 and LLaMA-2 show the potential of demographic information, user emotions and implicit feedback to improve task completion and the factual consistency of the generated responses. Our human evaluation shows the positive impact on user acceptance. For future work, we are planning to improve the quality of annotations in FEDI and to generalize our collection framework to other tasks and domains to increase the representativeness of FEDI. We are also planning to extend FEDI to be multimodal, including images that represent emotions, as well as audio signals.
9 Limitations
The training and validation dialogues in FEDI were synthetically generated using GPT-3.5-Turbo. Thus, there is a probability that some data is unfaithful, hallucinated, or even harmful Kumar et al. (2023); Zhang et al. (2023); Malaviya et al. (2023). In addition, some of these dialogues may seem artificial and unnatural due to potentially conflicting demographic information, e.g., language style contradicting with age or occupation, although we did not observe this in our analysis. In our experiments, we analyzed the toxicity of responses generated by the models using the Perspective API and found it to be negligible ($\leq 3\%$ for most cases). However, we acknowledge that the detector may not capture all the potential harmful content. The generated data may also contain positive stereotypes, i.e., seemingly-harmless words or patterns that are offensive to specific demographic groups, which are not marked by the detector Cheng et al. (2023).
Another limitation concerns the feedback scenarios represented in the feedback dialogues. Since they were generated artificially, it is possible that some user reactions appear to be unnatural, counterintuitive, and maybe not even addressing the underlying generation error. However, we did not observe these issues in our analysis.
Appendix A Task Descriptions
In the following, we provide details on the tasks included in FEDI and their slot values. Following Budzianowski et al. (2018), we distinguish requestable and informable slots, since this is necessary to calculate the task completion metrics in Section 7.
Post Office Services
FEDI includes dialogues from two basic services provided by post offices, customer support for parcel shipping and topping up a prepaid SIM card. In customer support for parcel shipping, the task is to help the user choose the right shipping box and delivery option for their needs (given the weight of the goods to be sent and the destination). Topping up a prepaid SIM card is less of an advisory service since customers usually know how much they want to recharge, their telephone number, and which telephone provider they are with. Table 6 lists the slots for each task.
Slot Name Informable Requestable Description
Parcel Shipping
Destination
The city and country of
destination; national or
international.
Weight
The weight of the item to be
shipped, lightweight (up to
5kg), average (up to 20kg),
heavy (up to 30kg).
Package Required
Whether or not a new
shipping box is required.
Delivery Option
Express or standard
delivery.
Country of Destination The destination country.
Shipping Box Name
Name of the best suitable
shipping box (small-sized,
medium-sized, large-sized),
based on the weight of the
item to be send.
Shipping Box Description
Brief description on why
the suggested shipping box
is a good choice.
Shipping Procedure
Description of the shipping
procedure (e.g., take the box
to the counter…)
Shipping Time
Expected delivery time, one
to three days for national,
four to six days for european,
and 3-4 weeks for inter-
national deliveries.
Top Up Prepaid SIM Card
Phone Number
Table or mobile phone
number with country code,
e.g., +39 XXX XXXXXXX
Phone Provider
The phone provider, e.g.
Vodafone, POSTE Mobile, …
Import Payment
The recharge amount, e.g.,
10 euro, 20 euro, 30 euro
Outcome Operation
If all required information
were provided, the system
asks the user to insert the
card for payment.
Request Ticket
Type of Service
The type of service for
which the user wants to
request support, i.e., Parcel
Choice or Recharge Phone.
Ticket Number
The ticket number generated
for the request.
Table 6: Slot values for parcel shipping and SIM card top-up
In modern post offices, service robots or other virtual agents are more commonly used to provide such services in a self-service manner. However, if something goes wrong, e.g., the shipping boxes are empty or the credit card was rejected, customers must have the option of requesting assistance from a human employee. In this case, the customer is asked to tell the agent the type of service they need assistance with. In turn, the agent creates a ticket for a human employee and returns the ticket number. We consider this as a kind of subtask to the other tasks. However, since the slots are the same, Table 6 lists this as Request Ticket.
Table 6 lists the annotations for each task. As intent annotations, we use the following values: parcel_choice, recharge_a_prepaid_phone and request_ticket.
Receptionist Services
For receptionist services, we only include one task in FEDI: access control. Table 7 shows the slots for this task.
Slot Name Informable Requestable Description
Building Access
Guest Name
The name of the person
who wants to access the
building.
Host Name
The name of the person
the guest wants to visit.
Host E-Mail
The E-Mail address of
the host.
Alternative
Host Name
An alternative host, e.g.,
in case the host is not
available today.
Alternative
Host E-Mail
E-Mail address of the
alternative host.
Meeting Date
and Time
Date and time of the
appointment.
Meeting Room
Identifier
Unique identifier of the
room where the meeting
will take place.
Verification
Call
The system can set up a
verification call to let the
host visually inspect the
guest and authorize
access.
Confirmation
to open Turn-
stile
This is a signal to the
system that controls the
turnstile to let the guest
enter.
Add. Safety
Information
Any additional safety
information, e.g., related
to COVID-19.
Table 7: Slot values for access control.
It is an essential task in hotels, office buildings, or other facilities with restricted access. Visitors usually need to register at the reception desk before being allowed to enter. As of today, electronic access controls (EAC) are more common than reception desks, especially in the case of office buildings, and they are becoming increasingly intelligent. In our case, we focus on a scenario in which a visitor has an appointment with an employee in an office building. To access the building, the visitor needs to provide the EAC with information about the appointment, e.g., the name of the host, date and time, and the room number. The EAC can then decide to grant access or to call the host for confirming the visitor’s identity. If necessary, the EAC can also provide additional safety information, e.g., hygiene guidelines. The intent annotation for utterances related to this task is building_access.
Customer Service in the Insurance Domain
For customer service in the insurance domain, we focus on question answering in the context of pet, health or heritage insurance, as we ll as bank transactions and account conditions. As a source, we use the insurance policies from POSTE Italiane, which are also available in English language [^16] POSTE Italiane Insurance Policies, last accessed 13. September 2023.. Table 8 lists the slots.
Slot Name Informable Requestable Description
Question Answering
Question
A question related to
one of the topics.
Type of Bills
If the user asks a question
regarding a specific pay-
ment slip, they need to
provide the type.
Evidence
The answer to the user’s
question.
Bill Form
Description
Description of the
specific payment form
(if the question was about
a payment form).
Bill Form
Name
Name of the payment
form (if the question was
about a payment form).
Bill Form
Payment
Procedure
Information on how to fill
the payment form (if the
question was about a pay-
ment form).
Table 8: Slot values for question answering.
In the past, customers called their insurance agent or visited their local bank branch for all questions related such topics. Today, it is more common to talk to chatbots or other service agents first and only in exceptional cases to human employees. Overall, we extracted 100 question-document pairs, i.e., questions paired with a paragraph that contains the answer, for bank transactions and account conditions, 78 for health, 84 for heritage, and 39 for pet insurance, from the POSTE documents. The intent annotation for utterances related to this task is question_answering.
Greeting
In the prompts for dialogue generation (see Appendix C), we instruct GPT-3.5 to have a separate turn at the beginning and ending of a dialogue in which both roles greet each other by also considering the generated background story. Such utterances have the intent annotation greeting. However, we do not consider this as a separate task in the sense of this work and do not evaluate it separately.
Appendix B Dataset Features
In this section, we provide additional details on the demographic information and the error and user reaction types used to create FEDI.
Demographic Information
We distinguish 12 different language styles, including Their Age and Job, Standard, Colloquial, Formal, Gutter, Polite, Informal, Regional Dialect, Social Dialect, Jargon, Slang, Age, five demogrphic cohorts, including Boomers (born between 1952 and 1962), Generation X (born between 1962 and 1977), Millenials (born between 1977 and 1992), Generation Z (born between 1992 and 2007), and Generation Alpha (born between 2007 and 2016), and an extensive variety of occupations (overall 1,155), sampled from The Gazette [^17] Available in GitHub (last accessed on 31 July 2023)., including among others jobs from the fields of science and technology, education, arts and entertainment, healthcare, or manufacturing. For each dialogue, we randomly sample a new value for each characteristics and apply simple plausibility checks, e.g., a person from Generation A can only be a pupil.
Error and User Reaction Types
To generate implicit user feedback, we use the error and user response type taxonomy proposed by Petrak et al. (2023). For generation errors in system utterances they define the following nine error types as relevant for task-oriented document-grounded dialogues:
  • Ignore Question — This error occurs when the system fails to address the user’s question. Instead of providing a relevant response or clarification, the system disregards the user’s input.
  • Ignore Request — A situation in which the system fails to take action on a user’s request. It can occur due to various reasons, such as misinterpretation of the request, technical limitations, or system glitches.
  • Ignore Expectation — This error happens when the system fails to fulfill the user’s expectations in terms of understanding and addressing their needs or requests accurately withing the context of the task.
  • Attribute Error — If the system fails to correctly extract or understand the necessary slots or attributes from the user’s utterance, this is called an attribute error.
  • Factually Incorrect — System responses that are factually wrong or inaccurate.
  • Topic Transition Error — A situation in which the system’s response abruptly shifts to a different or previously discussed topic without a logical connection or adequate context.
  • Conversationality — Bad conversationality occurs when the system fails to maintain a coherent and natural conversation flow, e.g., it repeats previous responses or contradicts itself without recognizing or asking for new or missing information.
  • Unclear Intention — This error is characterized by the robot’s failure to accurately address the user’s intended objective.
  • Lack of Sociality — If a system’s response doesn’t adhere to social conventions, fails to include basic greetings, or exhibit toxic and disrespectful behavior or language, this is called Lack of Sociality.
They also define an error type for common-sense errors, but found them to be rare in task-oriented document-grounded dialogues. For this reason, we do not consider this error type in our work.
For user reactions in response to such generation errors, they propose the following taxonomy:
  • Ignore and Continue — The user ignores the error and continues the conversation, e.g., "Okay. Let’s leave it like that.".
  • Repeat or Rephrase — Instead of ignoring the error in the system utterance, the user repeats or rephrases their original concern, e.g., "Actually, I wanted you to …".
  • Make Aware with Correction — The user makes the system aware of its error and provides a correction or response alternative, e.g., "Partly. This doesn’t take into account that …".
  • Make Aware without Correction — Instead of providing a correction or response alternative, the user just makes the system aware of its error, e.g., "You’re wrong.".
  • Ask for Clarification — In case of error, the user asks the system for clarification, e.g., "I’m not sure what you mean. Is it about …".
According to the authors, Repeat or Rephrase, Make Aware with Correction, and Ask for Clarification contain implicit user feedback.
Appendix C Prompts for Dialogue Generation and Annotation
In this section, we provide more details on the steps and prompts used for generating and annotating the dialogues of FEDI. Additionally added source data is highlighted in blue in the figures below.
JSON Schemes
As described in Section 4, we require GPT-3.5-Turbo to return all results in a predefined JSON scheme, which depends on the prompt, i.e., dialogue generation or annotation, and ensures that the returned values contain all required fields and is processable without human intervention. If the values returned by GPT-3.5 do not adhere to the required scheme, we drop the whole dialogue. Figure 6 shows an example for the annotation of emotions.
Figure 6: Instruction to return the results in json for emotion annotation.
We append these json schemes at the end of the prompts before sending them to GPT-3.5-Turbo. We basically provide the required fields and example values, and instruct it to return only utf-8 encoded characters and escape quotation marks (so that we can treat it as a string in Python). Please refer to our GitHub repository for all prompts and their json schemes1.
Feedback-Free Dialogues
For dialogue generation, we distinguish feedback-free and feedback dialogues. Figure 7 shows the instruction used to generate feedback-free dialogues.
Figure 7: Instruction for generating feedback-free dialogues.
For this, we provide GPT-3.5-Turbo with the demographic information, the role of the starting actor, and the task description. We require the model to use this information to generate a background story and to use this as an additional source for dialogue generation. We also instruct the model to return the utterance-level annotations for intents in this step.
Feedback Dialogues
The generation of feedback dialogues requires feedback scenarios as an additional source, which we generate in a separate step before dialogue generation using GPT-3.5-Turbo. We generate three feedback scenarios using the same promt for each new dialogue. Figure 8 shows the instruction.
Figure 8: Instruction for generating the feedback scenarios.
We provide the model with the task description and a list of three error and user reaction types. We then instruct the model to combine them uniquely and to generate three non-exclusive feedback scenarios that together form a story in the context of the task description. We also require the model to return a description and reasoning for each feedback scenario. We then use the returned feedback scenarios for the generation of feedback dialogues, which is shown in Figure 9.
Figure 9: Instruction for generating erroneous dialogues.
The instruction is longer and more detailed than the one used for generating the feedback-free dialogues (Figure 7). For example, it explicitely describes how to process error situations. Another difference is the length limitation. While feedback-free dialogues are restricted to 13 turns, we require feedback dialogues to have at least 13 turns. In practice, the length of the feedback dialogues is similar to the length of the feedback-free dialogues, but we observed that feedback dialogues are likely to be cut off without this requirement.
Continual Learning Dialogues
To create dialogues for continual learning, we create three additional versions of each feedback dialogue, each resolving one of the feedback scenarios from the original dialogue (until all feedback scenarios are solved). For this, we regenerate the system utterance with generation error. Figure 10 shows the instruction.
Figure 10: Instruction to regenerate the system utterance to replace the one with the generation error.
The instruction includes the dialogue up to the erroneous system utterance, the name of the task and the respective document (in the case of question answering). We mask the erroneous system utterance using <mask>, since we found this to produce replacements more coherent to the dialogue history. After replacing the erroneous system utterance, we regenerate its slot values. We also take advantage of the three-utterance-span constraint (see Section 4.2) and remove the following two utterances to ensure that the dialogue flow is not corrupted. This way, the conversation continues with the next regular user user utterance.
We experimented with different approaches for fixing errors in generated dialogues:
  • Using the implicit user feedback and the task description and instruct GPT-3.5-Turbo to rewrite the whole dialogue.
  • Providing GPT-3.5-Turbo with the whole dialogue and only instruct it to rewrite the affected turn.
  • Using the respective feedback scenario as additional input to regenerated the erroneous system utterance.
However, these approaches mostly resulted in inconsistent dialogues and off-topic, unnatural or incorrect system utterances. During prompt engineering, we found that the feedback itself is negligible for resolving feedback scenarios using our prompt from Figure 10, since it includes the original task description and requires the new system utterance to be generated in a friendly and polite manner.
Slot Annotations
Figure 11 shows our instruction used for generating slot values.
Figure 11: Instruction for slot annotation in a generated dialogue.
For this, we provide GPT-3.5-Turbo with the complete dialogue and distinguish between slots for each role (person and system). The slots to be annotated are provided in lists (including example values). We also instruct the model to just use sequences from the dialogue as slot values (to avoid hallucinated slot values).
Emotion Annotations
Figure 12 shows the instruction for emotion generation. We generate emotions just based on the dialogue context. We do not provide additional information, such as examples for each of the emotion. However, we additionally provide the number of utterances in the dialogue and those related to the user.
Figure 12: Instruction for generating emotions.
Appendix D Studies With Human Annotators
In this work, we conducted several studies that required human annotators: (1) the test data collection for FEDI, (2) the curation and assessment of the generated annotations and (3) the human evaluation of our trained models. While the human evaluation was conducted by one of the authors and an intern from our research group (who participated during his working hours), the test data collection and curation study was conducted with external participants who were hired for this purpose.
Application Criteria and Hiring Procedure
For participation, we required a formal application. Our criteria were as follows:
  • Enrollment in computational linguistics, linguistics, data and discourse studies, computer science, business informatics or comparable.
  • Fluent in reading, speaking and writing English.
  • Good communication and organization skills.
We considered a background in NLP, interest in conversational AI and experience in data annotation as a plus. We did not restrict the job advertisment to our university. Also, we did not take gender into consideration. We asked all applicants that fulfilled those criteria to take part in a recruitment test, in which we asked them to collect and annotate dialogues in a self-chat manner, given a random task description from our work. We then assessed and ranked their results based on (1) time needed for one dialogue, (2) annotation completness, (3) number of turns per dialogue, (4) avg. utterance length.
Overall, we received 11 applications, that all fulfilled our criteria. Eight of them passed the recruitment test and were hired for an hourly salary of 12,95\$. While all participants took part in the test data collection only two were involved in the data curation study.
Test Data Collection
For test data collection, we ransomly assigned participants to groups of two to collect the dialogues in one hour sessions dedicated to one task. For each task, we provided the task description, including slots with examples and four persona profiles (combinations of demographic information) and background stories as inspiration. However, we encouraged them to think about own persona profiles and background stories. For emotions, we provided lists of available options. For question answering, we also provided a set of at least 20 question-document pairs per topic.
For data collection, we use a self-developed web-based platform that allows to collect dialogues between two humans or between a human and a language generation model. Figure 13 shows the user interface.
Figure 13: The user interface of the data collection platform used to collect the test data.
[…]
Appendix E FEDI– Additional Analysis
In this section, we provide additional analysis about the composition of FEDI. Overall, FEDI consists of 8,852 dialogues, divided into 1,988 feedback-free and 6,864 feedback dialogues. Table 9 shows the distribution of dialogues in the dataset. Test for the feedback-free dialogues refers to the human-human collected test data.
Task
Feedback-Free
Dialogues
Feedback Dialogues
Version 1 Version 2 Version 3 Version 4
Train Dev Test Train Dev Train Dev Train Dev Train Dev
Parcel
Shipping
186 20 38 193 21 193 21 193 21 193 21
Top Up
SIM Card
187 20 39 193 21 193 21 193 21 193 21
Access
Control
183 20 42 215 23 215 23 215 23 215 23
Question
Answering
943 103 207 945 105 945 105 945 105 945 105
Per Split 1,499 163 326 1,546 170 1,546 170 1,546 170 1,546 170
Total 1,988 6,864
Table 9: Data splits included in FEDI and their sizes.
Demographic Information
Figure 14 shows the distribution of language styles, age ranges and occupations randomly sampled for background story generation.
Figure 14: The distribution of persona attributes represented in the background stories (excluding human-human test dialogues).
Language styles are almost equally weighted. In terms of occupations, the figure shows that jobs from the categories of business administration, service, industrial and manufacturing, and pupil largely outweigh the other categories, which makes sense in the context of the tasks and topics represented in FEDI  [^18] The original list did not provide categories. We generated them using GPT-3.5-Turbo.. Overall, we observe 693 unique job titles in FEDI. The figures do not show the distribution of names. We found 1,496 different names in the dialogues. 638 (42%) are unique, and 712 (47.59%) occur two to three times. The remaining 146 names occur four or more times throughout the entire dataset.
Emotions
The chart in Figure 15 show the distribution of emotions in the dialogues of FEDI (including error and test dialogues).
Figure 15: Illustration of the distribution of emotions in FEDI.
With 40.5%, Neutral is the most common emotion, followed by Curiosity (27.5%). Frustration and Confusion are relatively rare. We observe them mostly in the feedback dialogues. Others refers to emotions that are represented $\leq 5\%$, including Anger, Disgust, Fear, Surprise, and Stress.
Feedback Scenarios
Overall, we generated 4,714 feedback scenarios that are included in the 1,716 feedback dialogues of Version 1. Figure 16 shows the distribution of error and user reaction types.
Figure 16: Distribution of error and user reaction types in the feedback dialogues of FEDI.
Given that most of the dialogues are about question answering (Table 9), it is not surprising that Ignore Question is the most frequent error type. Table 10 shows the ten most commonly observed error and user reaction type combinations.
Error Type Feedback Type Frequency
1 Ignore Question Ignore and Continue 273
2 Ignore Request Ignore and Continue 208
3 Ignore Expectation Ignore and Continue 199
4 Unclear Intention Ask for Clarification 191
5 Ignore Question Repeat or Rephrase 187
6 Factually Incorrect
Make Aware With
Correction
166
7
Topic Transition
Error
Ask for Clarification 158
8 Attribute Error Repeat or Rephrase 156
9 Ignore Expectation Repeat or Rephrase 151
10 Lack of Sociality
Make Aware Without
Correction
141
Table 10: The table shows the most common error and user reaction type combinations included in FEDI.
Ignore Question and Ignore Request are two of the most frequent error types. While we observe the first one more common in question answering dialogues, the second one is more common in the other tasks. For both we observe that Ignore and Continue is the most frequent user reaction type, followed by Repeat or Rephrase. Unclear Intention is an error type mostly observed in parcel shipping, top up prepaid SIM card, and access control. The most frequently observed user reaction to this is Ask for Clarification. Based on absolute numbers, Factually Incorrect is the rarest error type. It’s mostly observed in question answering and in combination with Make Aware With Correction.
Appendix F Additional Details
In this section, we provide additional information on our experiments, including hyperparameters, input sequences and the results for our continual learning experiments.
Hyperparameters
For the experiments with feedback-free dialogues, we trained all models for five epochs, except for LLaMA-2 Touvron et al. (2023), since it took already five epochs to adapt the pretrained model to our prompting mechanism. For the experiment with feedback dialogues, we subsequently trained the best performing feedback-free models for ten epochs using the feedback data (ten epochs, since we have seen further improvements after the fifth epoch). We used a batch size of 32 and a learning rate of $55e-$ 5 with no warmup steps. As optimizer, we used the implementation of AdamW [^19] AdamW in the Pytorch documentation (last accessed 30. September 2023). Loshchilov and Hutter (2019) in Pytorch. Except for LLaMA-2, we fully-finetuned all models. For LLaMA-2, we only finetuned the LoRA Hu et al. (2022) weights, using a rank of 8, an alpha of 16, and a dropout rate of 0.05.
Input Sequences
Each model used in this work requires a different input sequence. In general, the components of the input sequence depend on the features used (e.g. user emotions or demographic information). Figure 17 shows the input sequence used for training and inference using FLAN-T5 Chung et al. (2022). Additionally added source data is highlighted in blue in the figures below.
Figure 17: Input sequence for FLAN-T5.
The target sequence includes the intent, slot values, and system response. It is basically the same as the last part of the input sequence for GPT-2 Radford et al. (2019), which is shown in Figure … (starting from <intent>, but without the special token).
Figure 18: Input sequence for GPT-2.
For inference with GPT-2, we use the same input sequence as for FLAN-T5 (Figure 17). For LLaMA-2, Figure 19 shows the input sequence.
Figure 19: Input sequence for LLaMA-2.
In contrast to FLAN-T5 and GPT-2, we use an instruction for training and inference with LLaMA-2. Similar to GPT-2, for inference, we only use the sequence up to the dialogue context.